(a) In a certain province,license plates consist of three letters (A-Z) and three digits (0-9). How many different plates can be issued? (b) If the province allows any six character mix (in any order) of 26 letters and 10 digits, how many unique plates are possible?
Thanks for the help

Answer by sudhanshu_kmr(1152)

a) number of possible ways = 26^3 * 10^3 = 17576000
b) number of possible ways = 36^6
( because for each place we can put any letter/digit out of total 36 i.e (10+26))

if any doubt, you are welcome to contact...